# OrbitZoo: Multi-Agent Reinforcement Learning Environment for Orbital Dynamics

### v1

@@ 1 Introduction
Since the dawn of the space age in 1957, humanity has successfully launched approximately 20,000 satellites into Earth’s orbit [18], of which only about 50% remain operational. These satellites are crucial for our daily life, providing critical services such as global communication, navigation, weather forecasting, Earth observation, and scientific research. However, these advances come with many problems.
@@
Earth’s orbital environment hosts an estimated 140 million debris objects, with approximately 1 million of these being larger than 1 cm – large and fast enough to cause catastrophic damage upon impact [17]. Collisions with space debris generate more debris, leading to further collisions. This chain-reaction known as the Kessler syndrome, results in an exponential increase in debris, threatening the long-term sustainability of Earth’s orbits [27]. If no further measures are taken to address this issue, Earth’s orbits could become unusable.
@@
More recently, the congestion in Lower Earth Orbit (LEO) is likely to undergo a major change of scale for mega-constellations of telecommunication satellites, leveraging tens of thousands of satellites [9]. The increasing density of satellites and debris, particularly in LEO [11], presents formidable challenges for Space Traffic Management (STM), related to the accurate monitoring and tracking of space objects and development of adequate decision-support frameworks [35], compounded by issues of data scarcity and uncertainty. Despite the growing availability of orbital data and tools, current solutions for satellite maneuvers – ranging from orbital transfers to collision avoidance – remain heavily reliant on manual processes. Human experts must navigate an ever-increasing complexity of scenarios, making critical decisions under time constraints with incomplete or inaccurate information [21]. As the volume of satellites and orbital debris continues rising, these traditional methods are quickly becoming unsustainable, which demands the development of new strategies to minimize our impact, such as debris removal [7] and faster, more capable, and autonomous intelligent systems for decision-making. Adaptive control systems [4] provide robustness to uncertainty but struggle with complex, dynamic, and nonlinear space systems. RL, on the other hand, excels in real-time adaptation to such environments [49].
@@
We explore how the application of RL to orbital dynamics is shaped by two primary factors and suffers from a lack of standardization. First, various tools are available to generate data, influencing the balance between simplicity and realism in the environment. Second, frameworks for developing RL-based missions within these environments are often built from scratch, calling for additional validation of the dynamics – a particularly challenging task in the field of orbital mechanics. To enable the effective application of RL to satellite maneuvering, we created OrbitZoo, an environment designed for both high-fidelity orbital data generation and RL development. The contributions of this paper are structured into three main categories:
@@
  • Data Generation: Built on Python and with a robust space dynamics library on its background, OrbitZoo generates high-fidelity orbital data by incorporating realistic forces and perturbations, providing accurate datasets essential for machine learning and strategy validation;
@@
  • Reinforcement Learning: OrbitZoo is standardized for RL research, leveraging the PettingZoo library [48] to support multi-agent reinforcement learning (MARL) with a Partially Observable Markov Decision Process (POMDP) structure. This integration enables the development, training, and benchmarking of intelligent satellite maneuvering strategies for single and multi-agent missions in cooperative, competitive, or mixed scenarios;
@@
  • Customizable Framework with Visualization: OrbitZoo’s modular design allows users to define scenarios, incorporate custom models, and adapt the environment to specific needs, with clear separation of abstraction levels. It also features an interactive visual component, making it versatile for a wide range of applications while providing an accessible way to understand orbital behaviors and decision-making processes.
@@ 2 Related Work
Existing efforts in this domain can be broadly categorized into four areas: high-fidelity tools for generating and simulating orbital dynamics data, reinforcement learning (RL) for orbital dynamics, multi-agent RL (MARL) environments, and efforts to bridge the reality gap.
@@
Orekit [33] is one of the most comprehensive open-source libraries for astrodynamics and orbital mechanics, offering advanced capabilities for precise orbit determination [34, 38, 14], propagation and associated uncertainties [2], attitude determination, and trajectory analysis [16]. Its modular, Java-based architecture allows users to model systems ranging from simple Newtonian attraction to highly complex scenarios incorporating detailed gravity fields and perturbations, such as atmospheric drag, solar radiation pressure, and third-body effects. While versatile, Orekit’s steep learning curve and reliance on Java can pose challenges for users unfamiliar with its technicalities.
Poliastro [39], a Python-native library, offers a more accessible alternative, featuring tools for orbit propagation and transfer planning. Although it lacks the high-fidelity modeling of Orekit, Poliastro integrates with the Cesium library [15] for accurate 3D geospatial visualization, making it suitable for simpler scenarios or users with limited expertise in orbital mechanics.
Systems Tool Kit (STK) [5], the leading commercial orbital simulation software, offers extensive capabilities for mission planning and operational analysis. Despite its advanced visualization features and high precision, STK’s substantial cost restricts its accessibility to well-funded organizations.
@@
Reinforcement learning has been widely applied to satellite maneuvering, leveraging custom-built environments and advanced dynamics models. Researchers have frequently used the Circular Restricted Three-Body Problem (CR3BP) model [36, 19, 46, 29] for simplified simulations, though it often fails to capture real-world dynamics and thrust perturbations, posing challenges for non-experts in orbital dynamics aiming to extend RL to realistic scenarios. For Low-Earth Orbit (LEO) station-keeping, [6] developed an RL framework using PPO with a 4th-order Yoshida integrator that accounted for gravitational forces and atmospheric drag. [13] explored orbital maneuvers in geostationary orbit (GEO), using A2C for a perigee raise. However, both studies relied on simplified dynamics, limiting their applicability to more complex scenarios. Orekit has also been central to RL environments tackling complex dynamics. For instance, [28] employed a DDPG-based framework for low-thrust transfers but lacked adaptability to multi-agent scenarios and dynamic uncertainties. Collision Avoidance Maneuvers (CAMs) have been another key focus. [26] introduced ColAvGym, a single-agent environment for CAM in LEO, leveraging Orekit for dynamics and PPO for learning. Using real Conjunction Data Messages (CDMs), the environment retroactively reconstructed satellite and debris orbits for training. Alternatively, [45] and [12] relied on synthetic methods to generate debris scenarios, using environments built on OpenAI’s Gym with Orekit and a Simplified General Perturbations model (SGP4), respectively. These studies explored various observation and reward designs, including temporal features captured through recurrent networks like DRQN, to handle the problem as a Partially Observable Markov Decision Process (POMDP).
Multi-Agent Reinforcement Learning Environments
@@
MARL has recently gained traction in orbital dynamics, with applications in satellite coordination and task assignment. [23] introduced the RL-Enabled Distributed Assignment (REDA) algorithm, using Poliastro and DQN for task allocation across satellite constellations. However, REDA lacked integration with realistic orbital data. Similarly, [53] applied Multi-Agent PPO (MAPPO) for multi-satellite observation planning, combining STK and MATLAB for custom simulation frameworks but without leveraging industry-standard astrodynamics libraries like Orekit.
Bridging the Reality Gap
A critical challenge in applying RL to orbital dynamics is bridging the Reality Gap between simulation and real-world operations. This includes incorporating high-fidelity dynamics, handling uncertainties, and validating simulation results against real-world data. For example, [26] developed ColAvGym, which used real Conjunction Data Messages (CDMs) to train agents for collision avoidance. [45] and [12] relied on synthetic data and simplified dynamics for training, but their environments did not fully capture the complexities of realistic orbital interactions.
@@
Work Multi-Agent RL Industry-Standard Simulator High-Fidelity Dynamics Continuous Control Realistic Bodies and Thrust Interactive Visualization Publicly Available
Kolosa (2019)
Miller (2019)
Armando (2020) ✓(drag only)
Federici (2021)
Sullivan (2021)
Casas (2022)
Bourriez (2023)
LaFarge (2023) ✓(third bodies only)
Zhang (2023)
Solomon (2024)
Holder (2024)
Kazemi (2024)
@@
As shown in Tab. 1, OrbitZoo stands out with multi-agent RL support, high-fidelity orbital dynamics, and integration with Orekit, an industry-standard simulator. Unlike prior works with simplified models, OrbitZoo incorporates perturbative forces, supports Cartesian and equinoctial representations, and offers realistic body dynamics, thrust modeling, and interactive visualization. Its public availability further establishes it as a comprehensive framework for advancing learning-based approaches in space operations.
3 Background: Challenges in Multi-Agent RL for Orbital Dynamics
Reinforcement Learning (RL) provides a framework for decision-making under uncertainty, where agents learn through trial and error to maximize cumulative rewards [47, 43]. In the context of multi-agent RL (MARL), multiple agents operate in a shared environment, each pursuing individual or collective goals. MARL presents unique challenges, particularly in orbital dynamics, where environments are partially observable, highly dynamic, and governed by complex physical interactions.
3.1 Theoretical Foundations of MARL
MARL builds upon Markov Decision Processes (MDPs) [47, 10] and Partially Observable Markov Decision Processes (POMDPs), commonly used to model scenarios where agents cannot access the full environment state [25]. A POMDP is defined by the tuple $\langle S,A,P,R,O,\Omega\rangle$. At each timestep $t$, an agent observes $o_{t}\in O$ based on the current state $s_{t}\in S$, selects an action $a_{t}\in A$, and transitions to a new state $s_{t+1}$ with probability $P(s_{t+1}|s_{t},a_{t})$, receiving a reward $R(s_{t},a_{t})$. The goal is to learn a policy $\pi(a_{t}|o_{t})$ that maximizes expected returns over time.
Deep RL algorithms such as DDPG [31] and PPO [42] have been instrumental in extending traditional RL to continuous control problems, albeit the topic had been approached before [44]. In MARL, the interaction between agents is often modeled as a multi-agent POMDP (MA-POMDP), where each agent $i$ observes $o^{i}_{t}$, performs an action $a^{i}_{t}$, and receives a reward $r^{i}_{t}$. Proximal Policy Optimization (PPO) and variants, such as Multi-Agent PPO (MAPPO) [52], are well-suited for these scenarios, allowing a stable decentralized or centralized training of agents in shared environments. This is particularly important for missions involving constellations of satellites, where each satellite may need to coordinate, compete, or cooperate with others to achieve shared objectives in complex – and potentially noisy – orbital dynamics.
@@ 3.2 Challenges in Bridging the Reality Gap
Simulating orbital dynamics introduces a significant reality gap due to the need for high-fidelity physics modeling and integration of real-world uncertainties. This section outlines key challenges encountered when applying RL to orbital dynamics. Basic concepts related to orbital mechanics are explored in 0.A.
Complex Orbital Dynamics
@@
Orbital dynamics are governed by Newton’s laws of motion and universal gravitation, requiring accurate modeling of perturbative forces, including atmospheric drag, solar radiation pressure, and third-body gravitational effects. Simplified propagators, such as the Simplified General Perturbations (SGP) model, are computationally efficient but fail to capture the precision required for RL tasks. As provided by Orekit [33], numerical propagators offer higher fidelity by incorporating complex perturbations, enabling realistic simulations for long-term trajectory predictions and mission-critical maneuvers.
Coordinate Systems and State Representations
Orbital states are traditionally represented using Keplerian or Cartesian coordinates. However, these systems suffer from singularities in special cases, such as circular or equatorial orbits. Equinoctial elements are preferred in RL applications as they avoid such singularities while maintaining the advantages of traditional representations.
Thrust Modeling and Control
@@
Thrust actions are typically modeled in local reference frames such as RSW (Radial, Along-track, and Cross-track). Polar parametrization offers a physically realistic and constrained action space, where the thrust is defined in magnitude, deviation angle, and azimuthal angle, providing greater flexibility for modeling satellite maneuvering capabilities. The transformation between polar and Cartesian thrust vectors ensures compatibility with existing numerical propagators.
Ephemeris Data and Validation
@@
Ephemerides provide time-stamped predictions of orbital positions and velocities, enabling the validation of simulated trajectories. For example, we show that Orekit’s numerical propagator performs well when validated against ephemeris data from Starlink satellites. By optimizing physical parameters, such as drag and reflection coefficients, the propagator closely approximates real-world satellite trajectories, helping to bridge the reality gap in RL applications.
3.3 Exploration in High-Dimensional Action Spaces
Exploration in continuous action spaces, such as those encountered in orbital dynamics, presents additional challenges. Traditional RL algorithms like DDPG rely on deterministic policies and Gaussian noise for exploration, which may be suboptimal in complex environments. Algorithms like PPO address this limitation by outputting stochastic policies, allowing agents to explore action spaces and learn optimal maneuvers adaptively.

### v2

@@ 1 Introduction
Since the dawn of the space age in 1957, humanity has successfully launched approximately 20,000 satellites into Earth’s orbit ESA2025SpaceDebris , of which only about 50% remain operational. These satellites are crucial for our daily life, providing critical services such as global communication, navigation, weather forecasting, Earth observation, and scientific research. However, these advances come with many problems.
@@
Earth’s orbital environment hosts an estimated 140 million debris objects, with approximately 1 million of these being larger than 1 cm – large and fast enough to cause catastrophic damage upon impact ESA2024EnvironmentReport . Collisions with space debris generate more debris, leading to further collisions. This chain-reaction known as the Kessler syndrome, results in an exponential increase in debris, threatening the long-term sustainability of Earth’s orbits kessler2010kessler . If no further measures are taken to address this issue, Earth’s orbits could become unusable.
@@
More recently, the congestion in Lower Earth Orbit (LEO) is likely to undergo a major change of scale for mega-constellations of telecommunication satellites, leveraging tens of thousands of satellites BERNHARD20231140 . The increasing density of satellites and debris, particularly in LEO boley , presents formidable challenges for Space Traffic Management (STM), related to the accurate monitoring and tracking of space objects and development of adequate decision-support frameworks MANFLETTI2023495 , compounded by issues of data scarcity and uncertainty. Despite the growing availability of orbital data and tools, current solutions for satellite maneuvers – ranging from orbital transfers to collision avoidance – remain heavily reliant on manual processes. Human experts must navigate an ever-increasing complexity of scenarios, making critical decisions under time constraints with incomplete or inaccurate information cream_esa . As the volume of satellites and orbital debris continues rising, these traditional methods are quickly becoming unsustainable, which demands the development of new strategies to minimize our impact, such as debris removal arshad and faster, more capable, and autonomous intelligent systems for decision-making. Adaptive control systems Annaswamy provide robustness to uncertainty but struggle with complex, dynamic, and nonlinear space systems. RL, on the other hand, excels in real-time adaptation to such environments TIPALDI20221 .
@@
We explore how the application of RL to orbital dynamics is shaped by two primary factors and suffers from a lack of standardization. First, various tools are available to generate data, influencing the balance between simplicity and realism in the environment. Second, frameworks for developing RL-based missions within these environments are often built from scratch, calling for additional validation of the dynamics – a particularly challenging task in the field of orbital mechanics. To enable the effective application of RL to satellite maneuvering, we created OrbitZoo, an environment designed for both fast high-fidelity orbital data generation and RL development. The contributions of this paper are structured into three main categories:
@@
  • Data Generation: Built on Python and with a robust space dynamics library on its background, OrbitZoo generates high-fidelity orbital data by incorporating realistic forces and perturbations, providing accurate datasets essential for machine learning and strategy validation, while making use of parallel computations for fast propagation;
@@
  • Reinforcement Learning: OrbitZoo is standardized for RL research, leveraging the PettingZoo library pettingzoo to support multi-agent RL (MARL) with a Partially Observable Markov Decision Process (POMDP) structure, while also making use of the parallelism provided by the space dynamics library to properly scale to systems with thousands of bodies. This integration enables the development, training, and benchmarking of intelligent satellite maneuvering strategies for single and multi-agent missions in cooperative, competitive, or mixed scenarios;
@@
  • Customizable Framework with Visualization: OrbitZoo’s modular design allows users to define scenarios with an arbitrary number of bodies with or without thrusting capabilities, incorporate custom models, and adapt the environment to specific needs, with clear separation of abstraction levels. It also features an interactive 3D visual component, making it versatile for a wide range of applications while providing an accessible way to understand orbital behaviors and decision-making processes.
@@ 2 Related Work
Existing efforts in this domain can be broadly categorized into four areas: high-fidelity tools for generating and simulating orbital dynamics data, RL for orbital dynamics, MARL environments, and efforts to bridge the reality gap.
@@
Orekit orekit is one of the most comprehensive open-source libraries for astrodynamics and orbital mechanics, offering advanced capabilities for precise orbit determination orekit_determination ; orekit_determination_2 ; orekit_determination_3 , propagation and associated uncertainties orekit_uncertainties , attitude determination, and trajectory analysis CucchiettiORBITAC . Its modular, Java-based architecture allows users to model systems ranging from simple Newtonian attraction to highly complex scenarios incorporating detailed gravity fields and perturbations, such as atmospheric drag, solar radiation pressure (SRP), and third-body effects. While versatile, Orekit’s steep learning curve and reliance on Java can pose challenges for users unfamiliar with its technicalities.
Poliastro poliastro , a Python-native library, offers a more accessible alternative, featuring tools for orbit propagation and transfer planning. Although it lacks the high-fidelity modeling of Orekit, Poliastro integrates with the Cesium library cesium for accurate 3D geospatial visualization, making it suitable for simpler scenarios or users with limited expertise in orbital mechanics. Systems Tool Kit (STK) STK , the leading commercial orbital simulation software, offers extensive capabilities for mission planning and operational analysis. Despite its advanced visualization features and high precision, STK’s substantial cost restricts its accessibility to well-funded organizations.
@@
RL has been widely applied to satellite maneuvering, often relying on custom-built environments and simplified dynamical models. Many studies use the Circular Restricted Three-Body Problem (CR3BP) miller ; federici ; sullivan ; lafarge ; bonasera for tractable simulations, though it falls short in capturing real-world perturbations, making it challenging for non-experts to extend RL to realistic space missions. For LEO station-keeping, herrera2020 implemented Proximal Policy Optimization (PPO) ppo with a 4th-order Yoshida integrator, modeling gravitational forces and atmospheric drag, while tafanidis used Soft Actor-Critic (SAC) sac with dynamics developed from scratch that also included third body forces and SRP.
Orekit has also been employed in RL frameworks with more accurate dynamics. Kolosa2019 , for instance, used Orekit and Deep Deterministic Policy Gradient (DDPG) ddpg for low-thrust transfers. In geostationary orbit (GEO), casas2022 used Orekit and a synchronous variant of A3C a2c for a perigee-raising maneuver. Despite these contributions, all relied on simplified or mission-specific physics without multi-agent or uncertainty support, limiting their generalization to more complex scenarios.
Collision Avoidance Maneuvers (CAMs) have received growing attention. kazemi_satellite_2024 introduced ColAvGym, a single-agent CAM environment in LEO using Orekit and PPO. It leverages real Conjunction Data Messages (CDMs) to retroactively reconstruct satellite and debris orbits for training. In contrast, Solomon2024Collision and bourriez2023 employed synthetic debris scenarios in Gym-based environments incorporating Orekit and an adaptation of SpaceNav spacenav . These studies investigated different observation and reward strategies, modeling the problem as a POMDP with temporal features captured by Long Short-Term Memory (LSTM) lstm layers in the Deep Q-Network (DQN) dqn architecture, showcasing how RL algorithms with discrete action spaces and recurrent networks can be useful in such scenarios.
Multi-Agent Reinforcement Learning Environments.
@@
MARL has recently gained traction in orbital dynamics, with applications in satellite coordination and task assignment. holder introduced the RL-Enabled Distributed Assignment (REDA) algorithm, using Poliastro and DQN for task allocation across satellite constellations. However, REDA lacked integration with realistic orbital data. Similarly, zhang applied Multi-Agent PPO (MAPPO) mappo for multi-satellite observation planning, combining STK and MATLAB for a custom simulation framework but without leveraging industry-standard astrodynamics libraries like Orekit. qingyu developed a MARL framework for cooperative formation flying of multiple spacecraft considering only Newtonian attraction, while dolan investigated multi-agent CAM using a framework nayak that aggregates local neighborhood information through a graph neural network (GNN) GNN , considering two dynamic models rao ; vallado2013fundamentals that did not account for SRP and third body forces.
3 Background: Challenges in Multi-Agent RL for Orbital Dynamics
A major challenge in applying RL to orbital dynamics lies in bridging the Reality Gap between simulation and real-world operations. This requires incorporating high-fidelity dynamics, accounting for uncertainties, and validating simulations against real-world data. According to what was discussed in Sec. 2, the essential capabilities of an orbital RL environment include: support for multi-agent interactions (cooperative and competitive), integration with industry-standard simulators such as Orekit or Poliastro (rather than custom-built models), high-fidelity dynamics with gravity fields and perturbations, support for algorithms with continuous control, realistic body and thrust modeling (capturing distinct spacecraft characteristics and continuous thrust integration), interactive real-time visualization, and public code availability.
@@ 3 Background: Challenges in Multi-Agent RL for Orbital Dynamics
Work
Multi-Agent
RL
Industry-Standard
Simulator
High-Fidelity
Dynamics
Continuous
Control
Realistic Bodies
and Thrust
Interactive
Visualization
Publicly
Available
Kolosa (2019) Kolosa2019
Miller (2019) miller
Herrera (2020) herrera2020 ✓(drag only)
Federici (2021) federici
Sullivan (2021) sullivan
Casas (2022) casas2022
Bonasera (2022) bonasera
Dolan (2023) dolan ✓(J2 only)
Bourriez (2023) bourriez2023
LaFarge (2023) lafarge ✓(third bodies only)
Zhang (2023) zhang
Qingyu (2023) qingyu
Holder (2024) holder
Solomon (2024) Solomon2024Collision
Kazemi (2024) kazemi_satellite_2024
@@
As shown in Tab. 1, OrbitZoo distinguishes itself by containing all the aforementioned capabilities. Unlike previous environments based on simplified models, OrbitZoo incorporates multiple perturbative forces, supports Cartesian, Keplerian, and equinoctial representations, enables realistic body and thrust modeling through diverse integration methods, propagates uncertainty, and offers several methods to assist in the development of RL missions or multi-agent systems analysis. Its open-source nature further establishes it as a comprehensive platform for advancing learning-based space operations.
3.1 Multi-Agent Reinforcement Learning
RL provides a framework for decision-making under uncertainty, where agents learn through trial and error to maximize cumulative rewards Sutton1998 ; silver . In the context of MARL, multiple agents operate in a shared environment, each pursuing individual or collective goals. MARL presents unique challenges in orbital dynamics, where environments are partially observable, highly dynamic, and governed by complex physical interactions.
MARL builds upon Markov Decision Processes (MDPs) Sutton1998 ; bertsekas2019reinforcement and Partially Observable Markov Decision Processes (POMDPs), commonly used to model scenarios where agents cannot access the full environment state kaelbling1998planning . A POMDP is defined by the tuple $\langle S,A,P,R,O,\Omega\rangle$. At each timestep $t$, an agent observes $o_{t}\in O$ based on the current state $s_{t}\in S$, selects an action $a_{t}\in A$, and transitions to a new state $s_{t+1}$ with probability $P(s_{t+1}|s_{t},a_{t})$, receiving a reward $R(s_{t},a_{t})$. The goal is to learn a policy $\pi(a_{t}|o_{t})$ that maximizes expected returns over time.
Deep RL algorithms such as DDPG ddpg and PPO ppo have been instrumental in extending traditional RL to continuous control problems, albeit the topic had been approached before smart2000practical . In MARL, the interaction between agents is often modeled as a multi-agent POMDP (MA-POMDP), where each agent $i$ observes $o^{i}_{t}$, performs an action $a^{i}_{t}$, and receives a reward $r^{i}_{t}$. PPO and variants, such as MAPPO mappo , are well-suited for these scenarios, allowing a stable decentralized or centralized training of agents in shared environments. This is particularly important for missions involving constellations of satellites, where each satellite may need to coordinate, compete, or cooperate with others to achieve shared objectives in complex – and potentially noisy – orbital dynamics.
@@ 3.2 Challenges in Bridging the Reality Gap
Simulating orbital dynamics presents a substantial reality gap, driven by the need for high-fidelity physical modeling and the incorporation of real-world uncertainties. This section highlights the main challenges of applying RL to orbital dynamics, whose understanding relies on the fundamental concepts presented in Appendix B.
Complex Orbital Dynamics.
@@
Orbital dynamics follow Newton’s laws of motion and gravitation, requiring accurate modeling of perturbative forces such as atmospheric drag, SRP, and third-body effects. Simplified propagators like the Simplified General Perturbations (SGP) model are computationally efficient but lack the precision required for RL tasks. In contrast, Orekit’s numerical propagators orekit provide high-fidelity simulations by incorporating complex perturbations, enabling realistic long-term trajectory prediction and maneuver planning. Notably, the choice of integration method in numerical propagation affects realism: classical fixed-step methods such as 4th-order Yoshida or RK4 runge-kutta are computationally efficient, while variable-step schemes like Dormand-Prince dormand-prince offer higher accuracy.
Coordinate Systems and State Representations.
Orbital states are commonly represented using Keplerian elements, but these suffer from singularities in cases like circular or equatorial orbits. Equinoctial elements avoid such issues and are often preferred in RL applications for their robustness. However, Cartesian coordinates remain essential not only for numerical integration but also for tasks like computing inter-body distances or collision probabilities. An RL framework should therefore support all major representations. In multi-agent settings, maintaining consistent reference frames and relative state representations is particularly challenging, as agents may operate in different orbital regimes.
Thrust Modeling and Control.
@@
Thrust actions are often defined in local reference frames like RSW (Radial, Along-track, Cross-track). A polar parametrization – using thrust magnitude, deviation angle, and azimuthal angle – provides a realistic and constrained action space, offering flexibility in modeling satellite maneuvers. This representation can be converted to Cartesian thrust vectors for compatibility with numerical propagators. In multi-agent settings, additional challenges arise as each spacecraft may have distinct thrust capabilities, fuel constraints, and control dynamics, making coordinated maneuver planning and learning more complex.
Exploration in High-Dimensional Action Spaces.
Exploration in continuous action spaces, such as those encountered in orbital dynamics, presents significant challenges, particularly in multi-agent settings where agents may operate in joint state and action spaces with cooperative or adversarial interactions. To address this complexity and the wide range of possible learning objectives, it is essential to design a modular framework that supports both on- and off-policy RL methods, as well as discrete and continuous action spaces. Such flexibility enables consistent experimentation and comparison across diverse mission scenarios. Moreover, having the capability to handle multi-objective RL (MORL) morl is important, as spacecraft missions often involve trade-offs (e.g., fuel vs. time vs. risk) that must be balanced within the policy learning process while coordinating multiple agents.
Ephemeris Data and Validation.
@@
Ephemerides offer time-stamped predictions of orbital positions and velocities, enabling validation of simulated trajectories. For instance, Orekit’s numerical propagator demonstrates high accuracy when compared with Starlink satellite ephemeris data. By tuning physical parameters – such as drag and reflection coefficients – the simulation closely matches real-world behavior, helping bridge the reality gap in RL applications.